OpenStudy (yato_gami):

I need help with this please! I want someone to go through the steps with me, not tell me the answer. Thanks~! ^_^

OpenStudy (yato_gami):

This is the question

OpenStudy (yato_gami):

Wait this is the wrong question xD lolol

OpenStudy (yato_gami):

There we go

OpenStudy (retireed):

I'm glad you said that because I was having a tough time trying to answer the question without giving the answers.

OpenStudy (yato_gami):

XD sorry about that lol

OpenStudy (retireed):

On the second submission, what do you want to know? I'm kind of confused.

OpenStudy (yato_gami):

Ah, wrong question again T-T shoot. They all look the same to me. Sorry! >.<

OpenStudy (yato_gami):

Ok, I'm 100% sure this is the question. Those last two aren't it. I just was to know the steps on how to solve it.

OpenStudy (retireed):

Open Study quit on me for a bit, but I got back in.

OpenStudy (yato_gami):

Oh ok

OpenStudy (retireed):

This is a good question.

OpenStudy (retireed):

The first equation is a linear function f(x) = 5x + 100 x is years (2) What is the value after 2 years?

OpenStudy (yato_gami):

So plug in 2 for x? Like this: f(2) = 5(2) + 100?

OpenStudy (retireed):

Yes.

OpenStudy (yato_gami):

Cool! So do we simplify that then?

OpenStudy (retireed):

What do you mean simplify? Just add them together. The second one is an exponential function.... g(x) = 100(1.04)^x again x = 2 What is the value of g(x)?

OpenStudy (yato_gami):

Wait so the first one we just add? How would we do that? If we're using order of operations, wouldn't 5(2) have to be solved first? And then add that too 100?

OpenStudy (retireed):

Yes you are correct I was combining two operations.

OpenStudy (yato_gami):

Ok so how would you do that? Can you go step by step?

OpenStudy (retireed):

How do you do what?

OpenStudy (retireed):

Since you are offline, I assume you are having internet issues. Here is what you need to do to solve the equations. Hopefully, it is enough explanation. f(2) = 5(2) + 100 g(2) = 100(1.04)^2 f(2) = 110 g(2) = 108.16 Then repeat the process for x = 20 f(20) = 5(20) +100 g(20) = 100(1.04)^20

OpenStudy (yato_gami):

Oh! Ok I see now! So you did f(2) = 5(2) + 100 f(2) = 10 + 100 f (2) = 110 right?

OpenStudy (yato_gami):

and for g(2) = 100(1.04)^2 it's g(2) = 100(1.0816) g(2) = 108.16 Thats how you did them both right? By using order of operations. @retirEEd

OpenStudy (yato_gami):

So then the account with the highest value in 2 years is 110, the first account yea? I'm going to try and solve for 20 years.

OpenStudy (retireed):

Keep going you are doing great.

OpenStudy (yato_gami):

:D okie! I'm solving for 20 years now! ^_^

OpenStudy (yato_gami):

Ok so the first account would be: f(20) = 5(20) + 100 f(20) = 100 + 100 f(20) = 200 And the second account would be: g(20) = 100(20)^2 g(20) = 100(400) g(20) = 40,000 So in 20 years, the second account would have the highest value. Right @retirEEd

OpenStudy (yato_gami):

Am I right? @retirEEd

OpenStudy (yato_gami):

Oh shoot xD I did the second one wrong I just noticed

OpenStudy (yato_gami):

Is the first one correct though?

OpenStudy (retireed):

Redo the second one. You beat me to it.

OpenStudy (yato_gami):

lolol okie!

OpenStudy (yato_gami):

Ok so for the second one I got: g(20) = 100(1.04)^20 g(20) = 100(2.191123143033419) g(20) = 219.11

OpenStudy (yato_gami):

It would still be the highest value but am I correct? @retirEEd

OpenStudy (retireed):

Yes g(20) is higher than f(20).

OpenStudy (yato_gami):

Ok thank you! ^_^